Rasool lent equal sums of money on simple interest for 4 years to Payal at 9% per annum and to Maihar at 8% per annum find the sum lent to each if the interest paid by Maihar is 600 rupees less than that paid by Payal
1
Expert's answer
2019-11-26T10:01:14-0500

Its a simple interest question,

Principal amount is the same, time is the same only total interest vary by Rs. 600.00

Interest rate paid by first case say R1=9%=0.09

Interest paid by second case say R2=8%=0.08

Total Interest paid by first case say I1

Total Interest paid by second case say I2=I1-600

Principal amount is P

Time is T=4 years

Formulae for total interest

"I=PRT"

Thus

"I\/R=PT"

For two different P & T both are same

Thus,

I1/R1=I2/R2

I1/0.09=(I1-600)/0.08

Thus I1=5400

and I2=5400-600=4800

By any case we can calculate principal amount

So through first case,

P=I/RT=5400/(0.09x4)=15000

Thus total amount paid by first case( i.e. by payal)=Principal amount+total interest=15000+5400=Rs. 20400

and

Thus total amount paid by second case( i.e. by maihar)=Principal amount+total interest=15000+4800=Rs.19800